# Quantum Complexity in Constrained Many-Body Models: Scars, Fragmentation, and Chaos

### v1

@@ III.3 Hilbert Space Fragmentation
Symmetries typically divide the Hilbert space into different subspaces that do not interact with each other and are characterized by conserved quantum numbers. However, due to the presence of kinetic constraints in the system, the Hilbert space may get “fragmented” into exponentially many different additional subspaces. This phenomenon is commonly known as Hilbert space fragmentation [20, 19, 22, 23]. The term was first coined in [18]. The number of subspaces due to symmetry normally stays constant with the number of lattice sites or at most grows polynomially. However, in the case of strong HSF, the behavior is opposite, and the number of subspaces can grow exponentially with system size $L$. The quantum states $|\psi\rangle$ in the dynamically disconnected spaces—commonly referred to as “Krylov subspaces”- remain confined to the same subspace under the action of the Hamiltonian. In other words, the full Hilbert space can be characterized by
$\mathcal{H}=\bigoplus_{i=1}^{\mathcal{N}}\mathcal{K}_{i}\ ,\ \ \mathcal{K}_{i}=\{|\psi\rangle,H|\psi\rangle,H^{2}|\psi\rangle,\cdots\}\ .$ (6)
@@
We observed that, for the symmetry-resolved subspace ($k=0$) of $H^{\mathcal{N}(0)}$, Hilbert space fragmentation is present. However, this is mainly dominated by trivial Krylov spaces each consisting of only a single basis state that gets annihilated by the Hamiltonian due to the kinetic constraint. Interestingly, as is known from the literature, the PPXPP model with the Rydberg constraint does not exhibit any fragmentation [41], which can be reverified from our analysis. In Fig. 10 in Appendix A, for $L=7$, we present a schematic of the fragmented subspaces obtained for both $H^{\mathcal{N}(0)}$ and the PPXPP model. We can see the additional condition coming from the Rydberg constraint eliminates all but the largest-dimensional Krylov space of $H^{\mathcal{N}(0)}$, thereby removing the fragmentation in the PPXPP model.
@@
Interestingly, we find that $H^{\mathcal{N}(1)}$ exhibits HSF with the number of fragmented subspaces increasing exponentially with the size of the system $L$, as shown in Fig. 2(a). Additionally, in this case, the ratio $d_{L}^{1}/\mathcal{D}_{L}^{1}$ approaches zero as $L$ approaches infinity, indicating that $H^{\mathcal{N}(1)}$ demonstrates a strong form of Hilbert space fragmentation (HSF), as depicted in Fig. 2(b). A schematic representation of the fragmented subspaces in zero momentum inversion symmetric sector for this model is shown in Fig. 1 for $L=10$. In contrast, for $H^{(\mathcal{N}(2)}$, although we observed HSF, the number of fragmented subspaces increases polynomially with the size of the system. In this case, the behavior of $H^{\mathcal{N}(3)}$ remains exactly the same as $H^{\mathcal{N}(1)}$. Notably, while the individual Hamiltonians $H^{\mathcal{N}(2)}$ and $H^{\mathcal{N}(3)}$ display distinct behaviors, their sum, in other words, $H^{\text{QGL}}$, shows no fragmentation at all. Apart from the trivial states that are annihilated by the Hamiltonian, the entire Hilbert space is instead highly connected. For $H^{\text{Pert}}_{\text{PPXPP}}$, which is a combination of the Hamiltonians $H^{\mathcal{N}(0)}$ and $H^{\mathcal{N}(1)}$, the HSF persists with dimension of the largest fragmented subspace higher than that obtained for $H^{\mathcal{N}(1)}$. We will discuss more about this observation in detail in Sec. V.
@@
In Table 2, we show the number of Krylov subspaces $(\mathcal{N}_{L})$, the dimension of the Hilbert space after the resolution of the symmetries $(\mathcal{D}_{L})$, and the dimension of the largest subspace $(d_{L})$ among them, for $H^{\mathcal{N}(1)}$, $H^{\mathcal{N}(2)}$ with increasing $L$. Here, we have also counted the trivial annihilated states.
Figure 2: Krylov subspaces and their scaling in fragmented model $H^{\mathcal{N}(1)}$. In (a) and (b), we illustrate how the number of Krylov subspaces, $\mathcal{N}_{L}$, and the ratio $d_{L}/\mathcal{D}_{L}$ behaves with system size $L$, respectively. The maximum system size considered is $L=22$. Both figures clearly indicate the presence of strong Hilbert space fragmentation (HSF). In contrast, $H^{\mathcal{N}(2)}$ (not shown here) displays a slower decay of $d_{L}/\mathcal{D}_{L}$ with increasing $L$.
@@
Figure 3: (a) Behavior of entanglement entropy for the eigenstates of kinetically constrained quantum many-body models $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$. In (a), we present the behavior of half-chain entanglement entropy ($\mathcal{S}_{L/2}$) across the part of the spectrum of $H^{\mathcal{N}(1)}$ belonging to the zero momentum and inversion symmetric sector for $L=18$. In contrast, for $H^{\mathcal{N}(2)}$, we obtained a similar behavior in (b) for the subspace that along with the translation symmetry ($k=0$ subspace) and inversion symmetry, has an additional spin-flip symmetry. In (c), we present the same plot as shown in (a), but highlighting the entanglement behavior of four specific Krylov subspaces using different colors. The three largest subspaces are of dimensions 749, 628, and 627, while the 50th largest subspace has a dimension of only 21. Notably, for $H^{\mathcal{N}(1)}$, the subspace with the largest dimension does not correspond to the highest entanglement entropy; instead, the highest entanglement comes from the second and third largest subspaces. Conversely, for $H^{\mathcal{N}(2)}$, as shown in (d), the largest Hilbert space, with dimension $d_{L}=1003$ (marked in blue), produces the highest entanglement entropy. The next two largest fragmented subspaces have dimensions of 965 and 711, marked in red and green, respectively.
Figure 4: Behavior of entanglement entropy for the eigenstates of the models QGL and the perturbed PPXPP with $\delta=0.09$. In (a) and (b) we plot the behavior of half-chain entanglement entropy ($\mathcal{S}_{L/2}$) for $H^{\text{QGL}}$ and $H^{\text{Pert}}_{\text{PPXPP}}$, respectively, for the eigenstates of the zero momentum and inversion symmetric sector for L=18.
@@ IV Quantum complexity study
As mentioned earlier, in our work, we carried out a thorough analysis of the interplay of quantum state preparation complexity measures, namely, entanglement and nonstabilizerness, and markers of quantum complexity such as quantum chaos. In the following, we will begin the discussion with the behavior of entanglement in these models and conduct a comparative study with other forms of complexities.
@@ IV.1 Entanglement
In the case of $H^{\mathcal{N}(2)}$, the dimensions of the three largest fragmented subspaces are 1003, 965, and 711, and the corresponding arches are marked in blue, red, and green, respectively. We can see from Fig. 3(d) that in this case, the arches follow the same ordering as their Hilbert space sizes. It is observed that in both the cases, $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, the Krylov subspaces with lower dimensions fail to exhibit thermal behavior since they occupy a very small portion of the full Hilbert space. Furthermore, these smaller subspaces tend to have low entanglement entropy, contributing to the variation along the y-axis in Figures 3(a) and (b). Therefore, our analysis characterizes these fragmented subspaces according to their entanglement-generating capabilities, revealing that the largest dimension space does not always correspond to the one that generates the highest amount of entanglement entropy.
Furthermore, we observe that as the HSF effect diminishes in the case of QGL, the symmetry-resolved subspace shows a consistent entanglement entropy (EE) profile as shown in Fig. 4(a). This results in a well-shaped arch, with very few eigenstates away from it. An interesting observation arises when we further consider the EE plot of $H^{\text{Pert}}_{\text{PPXPP}}$ in Fig. 4(b). In this case, the arch shape becomes significantly distorted. Rather, the maximum EE in this case increases almost linearly with energy.
IV.2 Chaotic behavior in the Model
@@
The entanglement entropy behavior, though, unveils the complexity associated with the eigenstates of the model, it does not provide a direct insight into the complexity of the physical models that we have discussed above. Specifically, it does not directly inform us about whether these models are integrable or non-integrable. For that purpose, we consider probing the chaotic behavior associated with the model through a systematic approach that we describe in detail below.
IV.2.1 Level Spacing Distribution
@@
We first study the energy level spacing distribution $P(s)$ where $s$ is the consecutive level spacing $s=e_{n+1}-e_{n}$ of the unfolded energy list. It is able to capture short range correlation in the spectrum. In our convention, $\{E_{n}\}$ is the original energy eigenvalue list and $\{e_{n}\}$ is the unfolded energy list. The unfolding is necessary to remove any kind of global effect in the spectrum due to the density of states of the system. This new energy set has an average level spacing one. The unfolding process is explained in the Appendix B. Chaotic models exhibit behavior similar to that of random matrices. In these models, the phenomenon of level repulsion causes the probability density function $P(s)$ to approach zero as the spacing $s$ approaches zero. As a result, the distribution conforms to the Wigner-Dyson distribution. In contrast, integrable models display a different behavior due to the uncorrelated nature of their energy eigenvalues, leading to level spacing that follows a Poisson distribution.
@@
In our case, we compute the level spacing distribution for all the models discussed above and systematically unveil the chaotic behavior. As a first step, it is essential to resolve all types of symmetries and constraints present in the model Hamiltonian. For example, in the case of $H^{\text{QGL}}$, the zero-momentum inversion-symmetric sector (with $\mathcal{I}=+1$) exhibits chaotic behavior, as shown in Fig. 5(a). However, for $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, merely resolving the symmetries is insufficient; one must also examine the fragmented subspaces of the Hilbert space separately. For instance, with $H^{\mathcal{N}(1)}$, in addition to focusing on the zero-momentum inversion symmetric space, we analyze the fragmented subspace with the largest dimension. The corresponding behavior is depicted in Fig. 5(b). Interestingly, for $H^{\mathcal{N}(2)}$, the additional spin-flip symmetry has a significant impact on the level spacing behavior. If we do not take this symmetry into account and consider the fragmented subspaces collectively, the level spacing distributions show characteristics consistent with a Poisson distribution. This can be seen in Fig. 12(a) in Appendix C. However, including either the spin-flip symmetry or fragmentation alone is insufficient, as it does not yield the exact Wigner-Dyson shape of the level spacing distribution. The hidden Wigner-Dyson (WD) distribution can only be revealed by carefully resolving this symmetry and independently examining the fragmented subspaces. In our analysis, we focus on the $\kappa=+1$ sector, and the behavior of the level spacing distribution obtained from the largest Krylov subspace is presented in Fig. 5(c).
@@
IV.2.2 Level spacing ratio

### v2

@@ III.3 Hilbert Space Fragmentation
Symmetries typically divide the Hilbert space into different subspaces that do not interact with each other and are characterized by conserved quantum numbers. However, due to the presence of kinetic constraints in the system, the Hilbert space may get “fragmented” into exponentially many different additional subspaces. This phenomenon is commonly known as Hilbert space fragmentation [21, 20, 23, 24]. The term was first coined in [19]. The number of subspaces due to symmetry normally stays constant with the number of lattice sites or at most grows polynomially. However, in the case of strong HSF, the behavior is opposite, and the number of subspaces can grow exponentially with system size $L$. The quantum states $|\psi\rangle$ in the dynamically disconnected spaces, commonly referred to as “Krylov subspaces”, remain confined to the same subspace under the action of the Hamiltonian. In other words, the full Hilbert space can be characterized by
$\mathcal{H}=\bigoplus_{i=1}^{\mathcal{N}}\mathcal{K}_{i}\ ,\ \ \mathcal{K}_{i}=\{|\psi_{i}\rangle,H|\psi_{i}\rangle,H^{2}|\psi_{i}\rangle,\cdots\}\ .$ (6)
@@
We observe that, in the symmetry-resolved subspace ($k=0$) of $H^{\mathcal{N}(0)}$, Hilbert space fragmentation is present. However, this fragmentation is predominantly trivial and is dominated by Krylov subspaces of dimension one. These correspond to basis states that are completely frozen by the kinetic constraint and are annihilated by the Hamiltonian, and therefore do not participate in any nontrivial dynamics. Interestingly, as is known from the literature, the PPXPP model with the Rydberg constraint does not exhibit any fragmentation [47], which can be reverified from our analysis. In Fig. 15 in Appendix A, for $L=7$, we present a schematic of the fragmented subspaces obtained for both $H^{\mathcal{N}(0)}$ and the PPXPP model. We can see the additional condition coming from the Rydberg constraint eliminates all but the largest-dimensional Krylov space of $H^{\mathcal{N}(0)}$, thereby removing the fragmentation in the PPXPP model.
@@
Interestingly, we find that $H^{\mathcal{N}(1)}$ exhibits HSF with the number of fragmented subspaces increasing exponentially with the size of the system $L$, as shown in Fig. 1(a). Additionally, in this case, the ratio $d_{L}^{1}/\mathcal{D}_{L}^{1}$ approaches zero as $L$ approaches infinity, indicating that $H^{\mathcal{N}(1)}$ demonstrates a strong form of Hilbert space fragmentation (HSF), as depicted in Fig. 1(b). Moreover, unlike $H^{\mathcal{N}(0)}$, the Krylov subspaces in this case consist of more than one basis state and are not trivially annihilated solely due to the kinetic constraint. Hence, it is thus important to distinguish between trivial and emergent Hilbert space fragmentation. In the case of $H^{\mathcal{N}(0)}$, as discussed earlier, the observed fragmentation originates from disconnected, frozen states rather than from dynamically generated invariant subspaces. Although the strong-HSF condition, $d_{L}/\mathcal{D}_{L}\to 0$ (see Fig. 16 in Appendix A) in the thermodynamic limit, is formally satisfied, this behavior is already visible at small system sizes. Consequently, the small value of $d_{L}/\mathcal{D}_{L}$ does not indicate emergent or dynamical fragmentation, but instead reflects the abundance of trivially disconnected states.
@@
A schematic representation of the fragmented subspaces of $H^{\mathcal{N}(1)}$ in zero momentum inversion symmetric sector for this model is shown in Fig. 2 for $L=10$. In contrast, for $H^{\mathcal{N}(2)}$, although we observed HSF, the number of fragmented subspaces increases polynomially with the size of the system. The fragmentation structure of $H^{\mathcal{N}(3)}$ remains exactly the same as $H^{\mathcal{N}(1)}$ since both are related by the spin-flip symmetry. Notably, while the individual Hamiltonians $H^{\mathcal{N}(2)}$ and $H^{\mathcal{N}(3)}$ display distinct behaviors, their sum, in other words, $H^{\text{QGL}}$, shows no fragmentation at all. Apart from the trivial states that are annihilated by the Hamiltonian, the entire Hilbert space is instead highly connected.
@@
In Table 2, we show the number of Krylov subspaces $(\mathcal{N}_{L})$, the dimension of the Hilbert space after the resolution of the symmetries $(\mathcal{D}_{L})$, and the dimension of the largest subspace $(d_{L})$ among them, for $H^{\mathcal{N}(1)}$, $H^{\mathcal{N}(2)}$ with increasing $L$. Here, we have also counted the trivial annihilated states. For $H^{\text{Pert}}_{\text{PPXPP}}$, which is a combination of the Hamiltonians $H^{\mathcal{N}(0)}$ and $H^{\mathcal{N}(1)}$ ( $H^{\mathcal{N}(2)}$), the HSF persists with richer structure of fragmented subspaces than that obtained for $H^{\mathcal{N}(1)}$ ($H^{\mathcal{N}(2)}$) (as can be seen from Table 3). We will discuss more about this observation in detail in Sec. V and show their comparisons with $H^{\mathcal{N}(0)}$, along with $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$ (Fig. 16) in Appendix B.
@@
Figure 3: Behavior of entanglement entropy for the eigenstates of kinetically constrained quantum many-body models $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$. In (a), we present the behavior of half-chain entanglement entropy ($S_{L/2}$) for the part of the spectrum of $H^{\mathcal{N}(1)}$ belonging to the zero momentum and inversion symmetric sector for $L=18$. For $H^{\mathcal{N}(2)}$, we obtained a similar behavior in (b) for the subspace that along with the translation symmetry ($k=0$ subspace) and inversion symmetry, has an additional spin-flip symmetry. In (c), we present the same plot as shown in (a), but highlighting the entanglement behavior of four specific Krylov subspaces using different colors. The three largest subspaces are of dimensions 749, 628, and 627, while the 50th largest subhas a dimension of only 21. Notably, for $H^{\mathcal{N}(1)}$, the subspace with the largest dimension does not correspond to the highest entanglement entropy; instead, the highest entanglement comes from the second and third largest subspaces. Conversely, for $H^{\mathcal{N}(2)}$, as shown in (d), the largest Hilbert space, with dimension $d_{L}=1003$ (marked in blue), produces the highest entanglement entropy. The next two largest fragmented subspaces have dimensions of 965 and 711, marked in red and green, respectively.
@@ IV Quantum complexity study
As mentioned earlier, in our work, we carried out a thorough analysis of the interplay of quantum state preparation complexity measures, namely, entanglement and nonstabilizerness, and markers of dynamical quantum complexity such as quantum chaos. In the following, we will begin the discussion with the behavior of entanglement in these models and conduct a comparative study with other forms of complexities.
@@ IV.1 Entanglement
Figure 4: Behavior of entanglement entropy for the eigenstates of the models QGL. Here we plot the behavior of half-chain entanglement entropy ($S_{L/2}$) of the eigenstates of $H^{\text{QGL}}$ in the zero momentum and inversion symmetric sector obtained for $L=18$. The behavior shows smooth variations of bipartite entanglement with energy.
@@
In the case of $H^{\mathcal{N}(2)}$, the dimensions of the three largest fragmented subspaces are 1003, 965, and 711, and the corresponding arches are marked in blue, red, and green, respectively. We can see from Fig. 3(d) that in this case, the arches follow the same ordering as their Hilbert space sizes. It is observed that in both the cases, $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, the Krylov subspaces with lower dimensions fail to exhibit thermal behavior since they occupy a very small portion of the full Hilbert space. Furthermore, these smaller subspaces tend to have low entanglement entropy, contributing to the variation along the y-axis in Figures 3(a) and (b). Therefore, our analysis characterizes these fragmented subspaces according to their entanglement-generating capabilities, revealing that the largest dimension space does not always correspond to the one that generates the highest amount of entanglement entropy. Thereafter, we observe that as the HSF effect diminishes in the case of QGL, the symmetry-resolved subspace shows a consistent entanglement entropy (EE) profile as shown in Fig. 4. This results in a well-shaped arch, with very few eigenstates away from it.
Figure 5: Growth of the entanglement entropy following a quench from an initial product state $|\psi_{i}\rangle$, chosen to belong to a specific Krylov subspace of the Hamiltonian, under (a) $H^{\mathcal{N}(1)}$ and (b) $H^{\mathcal{N}(2)}$. The long-time saturation value of the entanglement entropy increases with the dimension of the Krylov subspace accessed during the dynamics. Results are shown for $L=18$.
A qualitatively different picture emerges when we focus on dynamical entanglement generation. Here, we prepare an initial product state $|\psi_{i}\rangle$ belonging to a specific Krylov subspace and study its unitary time evolution under $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$. In contrast to the static eigenstate analysis, the long-time saturation value of the entanglement entropy is found to correlate strongly with the dimension of the Krylov subspace explored during the dynamics. As shown in Fig. 5, quenches within higher-dimensional Krylov subspaces lead to systematically larger entanglement entropy at long times, even though such a direct correspondence between Krylov space dimension and maximum entanglement is absent for $H^{\mathcal{N}(1)}$ at the level of individual eigenstates (revisit Fig. 3(c)).
IV.2 Stabilizer Rényi Entropy
The next aspect of quantum state preparation complexity discussed in our work is the nonstabilizerness. A useful way to quantify nonstabilizerness is by measuring how far a quantum state is from being a stabilizer state, which is referred to as the Stabilizer Rényi Entropy (SRE). For a pure state $|\Psi\rangle$ of $L$ qubits, the SRE is defined as [49]:
In our study, we examine a maximum system size of $L=12$ and plot the SRE behavior for all the considered models in Fig. 6. Similar to the behavior of entanglement as illustrated in Figs. 3 and 4, the SRE value obtained for the symmetry resolved fragmented sectors of $H^{\mathcal{N}(1)}$ (Fig. 6(a)) and $H^{\mathcal{N}(2)}$ (Fig. 6(b)) shows a more spread behavior than the Quantum Game of Life (Fig. 7), which shows smooth behavior and remains nearly flat with variations in energy. This flatness indicates that the eigenstates in the middle of the spectrum exhibit some variation in terms of entanglement while sharing the same SRE content. This behavior resembles the scenario where quantum states remain maximally localized in the Fock basis, resulting in flat Shannon information entropy as discussed in [57]. In case of $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, we separately plot the SRE behavior for a few of the largest subspaces to compare the SRE generating power of the dynamically disconnected Krylov sectors. Figure 6(c) reveals that for $H^{\mathcal{N}(1)}$, the highest and the second-highest dimensional subspaces generate almost the same SRE value. This is different from the entanglement behavior we observed in Fig. 3(c). Whereas, for $H^{\mathcal{N}(2)}$, Fig. 6(d) shows many of the eigenstates of the second-highest subspace from the middle of the spectrum show higher SRE values than those obtained for the highest-dimensional sector.
Figure 7: Behavior of the stabilizer Rényi entropy ($\mathcal{M}$) for eigenstates of the QGL model. We plot $\mathcal{M}$ for the eigenstates of $H^{\text{QGL}}$ in the zero-momentum, inversion-symmetric sector for $L=12$. For the bulk eigenstates the stabilizer Rényi entropy remains nearly independent of energy across the spectrum, indicating a uniform degree of non-stabilizerness among the eigenstates.
However, in quenched dynamics, the saturation value of the SRE, similar to the entanglement entropy, is strongly correlated with the Krylov-subspace dimension for $H^{\mathcal{N}(1)}$, while for $H^{\mathcal{N}(2)}$ the saturation of the SRE becomes insensitive to the increase in the Krylov dimension, with the three largest subspaces producing nearly identical values as depicted in Fig. 8.
Our analysis therefore reveals that both the dynamical and static behaviors of these rule-based models exhibit a rich quantum-state–preparation complexity, as probed through entanglement and non-stabilizerness. In particular, for the HSF model, these complementary diagnostics can distinguish the underlying dynamics across different regimes, reflecting the distinct capacities of dynamically disconnected sectors to generate quantum resources.
Figure 8: Growth of stabilizer Rényi entropy following a quench from an initial product stabilizer state $|\psi_{i}\rangle$, chosen to belong to a specific Krylov subspace of the Hamiltonian, under (a) $H^{\mathcal{N}(1)}$ and (b) $H^{\mathcal{N}(2)}$. For $H^{\mathcal{N}(1)}$, the long-time saturation value of SRE increases with the dimension of the Krylov subspace accessed during the dynamics. However, for $H^{\mathcal{N}(2)}$, the first three highest Krylov spaces saturate to almost same SRE value. Here we consider $L=12$.
IV.3 Chaotic behavior in the Model
@@
The behavior of entanglement entropy and SRE, though, unveils the complexity associated with the eigenstates of the model, it does not provide a direct insight into the complexity of the physical models that we have discussed above. Specifically, it does not directly inform us about whether these models are integrable or non-integrable. For that purpose, we consider probing the chaotic behavior associated with the model through a systematic approach that we describe in detail below.
IV.3.1 Level Spacing Distribution
@@
We first study the energy level spacing distribution $P(s)$ where $s$ is the consecutive level spacing $s=e_{n+1}-e_{n}$ of the unfolded energy list. It is able to capture short range correlation in the spectrum. In our convention, $\{E_{n}\}$ is the original energy eigenvalue list and $\{e_{n}\}$ is the unfolded energy list. The unfolding is necessary to remove any kind of global effect in the spectrum due to the density of states of the system. This new energy set has an average level spacing one. The unfolding process is explained in the Appendix C. Chaotic models exhibit behavior similar to that of random matrices. In these models, the phenomenon of level repulsion causes the probability density function $P(s)$ to approach zero as the spacing $s$ approaches zero. As a result, the distribution conforms to the Wigner-Dyson distribution. In contrast, integrable models display a different behavior due to the uncorrelated nature of their energy eigenvalues, leading to level spacing that follows a Poisson distribution.
@@
In our case, we compute the level spacing distribution for all the models discussed above and systematically unveil the chaotic behavior. As a first step, it is essential to resolve all types of symmetries and constraints present in the model Hamiltonian. For example, in the case of $H^{\text{QGL}}$, the zero-momentum inversion-symmetric sector (with $\mathcal{I}=+1$) exhibits chaotic behavior, as shown in Fig. 9(a). However, for $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, merely resolving the symmetries is insufficient; one must also examine the fragmented subspaces of the Hilbert space separately. For instance, with $H^{\mathcal{N}(1)}$, in addition to focusing on the zero-momentum inversion symmetric space, we analyze the fragmented subspace with the largest dimension. The corresponding behavior is depicted in Fig. 9(b). Interestingly, for $H^{\mathcal{N}(2)}$, the additional spin-flip symmetry has a significant impact on the level spacing behavior. If we do not take this symmetry into account and consider the fragmented subspaces collectively, the level spacing distributions show characteristics consistent with a Poisson distribution. This can be seen in Fig. 18(a) in Appendix D. However, including either the spin-flip symmetry or fragmentation alone is insufficient, as it does not yield the exact Wigner-Dyson shape of the level spacing distribution. The hidden Wigner-Dyson (WD) distribution can only be revealed by carefully resolving this symmetry and independently examining the fragmented subspaces. In our analysis, we focus on the $\kappa=+1$ sector, and the behavior of the level spacing distribution obtained from the largest Krylov subspace is presented in Fig. 9(c).
@@
IV.3.2 Level spacing ratio